OpenStudy (anonymous):

What is 15% off of $75

OpenStudy (anonymous):

$63.75

OpenStudy (anonymous):

75*0.15=11.25 75-11.25=63.75

OpenStudy (anonymous):

75x0.15 =11.25 75-11.25 = 63.75

OpenStudy (anonymous):

You can set up a ratio for this type of problem \[\frac{15}{100} = \frac{x}{75}\] and solve for x

OpenStudy (anonymous):

that will find 15% of 75, if you need to show your work. or use a calculator. do what @OrangeMaster did on step two, and subtract that 15% you found from 75 to find the discount